During its first year of operations, Ayayai Corp.had these transactions pertaining to its common st tions, Ayayai Corp. had these transactions pertaining to its common stock Jan. 10 Issued 25,000 shares for cash at $ 4 per share July 1 Issued 50,000 shares for cash at $7 per share al journalize the transactions, assuming that the common stock has a par value of $ 4 per share b) Journalize the transactions, assuming that the common stock is no-par with a stated value of $ 1 per share Record journal entries in the order presented in the problem Credit occount ites are automaticolly Indlented when amount is entered, De not indent manually itontry ls equired,select No Entry for the account and enter O for the amounts)
Debit Credit No. Date Account Titles and Explanation
Answer #1
Debit Credit
a Jan. 10 Cash 100000 =25000*4
       Common Stock 100000
July 1 Cash 350000 =50000*7
       Common Stock 200000 =50000*4
       Paid in capital in excess of par value-Common Stock 150000
b Jan. 10 Cash 100000 =25000*4
       Common Stock 25000 =25000*1
       Paid in capital in excess of stated value-Common Stock 75000
July 1 Cash 350000 =50000*7
       Common Stock 50000 =50000*1
       Paid in capital in excess of stated value-Common Stock 300000